A card is selected at random from a standard deck of playing cards.
Compute the probability that the card is a face card (jack, queen, or king). (Enter your probability as a fraction.)

Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

It is 3/13.
There are 4 jacks, 4 queens, 4 kings, or 12 face cards
There are 52 cards
12/52=3/13

Answer by stanbon(75887)

A card is selected at random from a standard deck of playing cards.
Compute the probability that the card is a face card (jack, queen, or king). (Enter your probability as a fraction.)
---
# of ways to succeed:: 3*4 = 12 face cards
# of possible choices:: 52 cards
----------------
P(select a face card) = 12/52 = 3/13
--------------
Cheers,
Stan H.
----------